# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ Appendix F Lagrangian Lower Bound for Gaussians: Proof of Theorem 4
$T_{\hat{\mathcal{F}},r}^{-1}({\boldsymbol{\mu}})=\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\frac{1}{2\sigma^{2}}\frac{(r-{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}))^{2}}{\|\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\right\}$
@@ F.1 Bounds on Sample complexity: Proof of Corollary 1 Part (a)
Part i. Then, we get $\frac{2\sigma^{2}K}{C_{\mathrm{known}}}(1+\mathfrak{s}_{\tilde{\mathbf{A}}})\leq T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\leq\frac{2\sigma^{2}K}{C_{\mathrm{known}}}$, where $C_{\mathrm{known}}=\min_{\boldsymbol{\pi}^{\prime\prime}\in\nu_{\mathcal{F}}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})}d_{\boldsymbol{\pi}^{\prime\prime}}^{2}$.
Part ii. Additionally, $T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\geq\frac{H}{\kappa^{2}}(1+\mathfrak{s}_{\tilde{\mathbf{A}}})$, where $H$ is inversely proportional to the sum of squares of gaps and $\kappa_{\mathrm{known}}$ is the condition number of a sub-matrix of $\mathbf{A}$ consisting ${K}$ linearly independent active constraints for $\boldsymbol{\pi}^{\star}$.
@@
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})-\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})$ $=\min_{\boldsymbol{l}\in\mathcal{L}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}})}\left\{\frac{1}{2\sigma^{2}}\frac{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\right\}$
@@
Let us remind, due to pessimistic choice of $\tilde{\mathbf{A}}$, $\mathcal{F}\subseteq\hat{\mathcal{F}}$. $\boldsymbol{\pi}^{\prime}$ is a neighbor of $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}$ if it is an extreme point in the polytope $\hat{\mathcal{F}}$ and shares (K-1) active constraints with $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}$. Then $\boldsymbol{\pi}^{\star}_{\mathcal{F}}$ and $\boldsymbol{\pi}^{\prime\prime}$ lies in the interior of $\hat{\mathcal{F}}$ i.e, they can be expressed as a convex combination of $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}$ and $\boldsymbol{\pi}^{\prime}$. Let, $\exists 0\leq t_{1}\leq 1:\boldsymbol{\pi}^{\star}_{\mathcal{F}}=t_{1}\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}+(1-t_{1})\boldsymbol{\pi}^{\prime}$ and $\exists 0\leq t_{1}\leq 1:\boldsymbol{\pi}^{\prime\prime}=t_{2}\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}+(1-t_{2})\boldsymbol{\pi}^{\prime}$ Then, $(\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}^{\prime})=t_{1}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime})$ and $(\boldsymbol{\pi}^{\prime\prime}-\boldsymbol{\pi}^{\prime})=t_{2}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime})$. Then,
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})-\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$
$=\min_{\boldsymbol{\pi}^{\prime\prime}\in\nu_{\mathcal{F}}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})}\frac{1}{2\sigma^{2}}\frac{\|\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}^{\prime\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}^{\prime\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\min_{\boldsymbol{l}\in\mathcal{L}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}})}\left\{\frac{1}{2\sigma^{2}}\frac{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\right\}$
$=\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}})}\frac{1}{2\sigma^{2}}\frac{(t_{1}-t_{2})^{2}\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{(t_{1}-t_{2})^{2}\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\min_{\boldsymbol{l}\in\mathcal{L}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}})}\left\{\frac{1}{2\sigma^{2}}\frac{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\right\}$
$=\min_{\boldsymbol{l}\in\mathcal{L}}\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}$
$\leq\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\mathfrak{s}_{\tilde{\mathbf{A}}}$ (14)
@@
Part I. Since $\min_{\boldsymbol{l}\in\mathcal{L}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\geq 0$ by the design of the Lagrangian and the constraints, we get
@@
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\left(1+\mathfrak{s}_{\tilde{\mathbf{A}}}\right)\leq\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})$
$\implies$ $T_{\hat{\mathcal{F}},0}^{-1}({\boldsymbol{\mu}})\left(1+\mathfrak{s}_{\tilde{\mathbf{A}}}\right)\leq T_{\mathcal{F},0}^{-1}({\boldsymbol{\mu}})$ (16)
@@
$T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\geq\left(1+\mathfrak{s}_{\tilde{\mathbf{A}}}\right)\frac{2\sigma^{2}}{\min_{\boldsymbol{\pi}^{\prime\prime}\in\nu_{\mathcal{F}}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})}d_{\boldsymbol{\pi}^{\prime\prime}}^{2}}$
@@ F.2 Impact of unknown linear constraints: Proof of Corollary 1 Part (b)
Part (b) Let $d_{\boldsymbol{\pi}}^{2}\mathrel{\triangleq}\frac{\|\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}\|_{2}^{2}}$ be the norm of the projection of ${\boldsymbol{\mu}}$ on the policy gap $(\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi})$. Then, the characteristic time $T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})$ satisfies $T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\geq\frac{H}{\kappa^{2}}\left(1-\frac{\epsilon}{\gamma}\right)$. $H$ is the sum of squares of gaps. $\kappa$ is the condition numbers of a sub-matrix of $\mathbf{A}$ that consists ${K}$ linearly independent active constraints for $\boldsymbol{\pi}^{\star}$.
@@
Here, $H=\frac{2\sigma^{2}}{\Delta^{2}}$, where $\Delta^{2}=\sum_{a=1}^{K}({\boldsymbol{\mu}}^{*}-{\boldsymbol{\mu}}_{a})^{2}$, i.e the sum of squares of sub-optimal gaps in the arms and $\kappa$ is the condition number of a sub-matrix of $\mathbf{A}$ consisting at least ${K}$ linearly independent active constraints for $\boldsymbol{\pi}^{\star}$. Referring to the proof structure of Corollary 2 in  (Carlsson et al.,, 2024) is of independent interest.
@@
$T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\geq\frac{H}{\kappa^{2}}\left(1+\mathfrak{s}_{\tilde{\mathbf{A}}}\right)$

### v2

@@ Appendix F Lagrangian Lower Bound for Gaussians: Proof of Theorem 4
$T_{\hat{\mathcal{F}},r}^{-1}({\boldsymbol{\mu}})=\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\frac{1}{2\sigma^{2}}\frac{(r-{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}))^{2}}{\|\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\right\}$
@@ F.1 Bounds on Sample complexity: Proof of Corollary 1 Part (a)
Part i. Then, we get $\frac{2\sigma^{2}K}{C_{\mathrm{known}}}(1+\mathfrak{s}_{\tilde{A}})\leq T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\leq\frac{2\sigma^{2}K}{C_{\mathrm{known}}}$, where $C_{\mathrm{known}}=\min_{\boldsymbol{\pi}^{\prime\prime}\in\nu_{\mathcal{F}}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})}d_{\boldsymbol{\pi}^{\prime\prime}}^{2}$.
[…]
@@
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})-\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})$ $=\min_{\boldsymbol{l}\in\mathcal{L}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}})}\left\{\frac{1}{2\sigma^{2}}\frac{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\right\}$
@@
Let us remind, due to pessimistic choice of $\tilde{A}$, $\mathcal{F}\subseteq\hat{\mathcal{F}}$. $\boldsymbol{\pi}^{\prime}$ is a neighbor of $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}$ if it is an extreme point in the polytope $\hat{\mathcal{F}}$ and shares (K-1) active constraints with $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}$. Then $\boldsymbol{\pi}^{\star}_{\mathcal{F}}$ and $\boldsymbol{\pi}^{\prime\prime}$ lies in the interior of $\hat{\mathcal{F}}$ i.e, they can be expressed as a convex combination of $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}$ and $\boldsymbol{\pi}^{\prime}$. Let, $\exists 0\leq t_{1}\leq 1:\boldsymbol{\pi}^{\star}_{\mathcal{F}}=t_{1}\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}+(1-t_{1})\boldsymbol{\pi}^{\prime}$ and $\exists 0\leq t_{1}\leq 1:\boldsymbol{\pi}^{\prime\prime}=t_{2}\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}+(1-t_{2})\boldsymbol{\pi}^{\prime}$ Then, $(\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}^{\prime})=t_{1}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime})$ and $(\boldsymbol{\pi}^{\prime\prime}-\boldsymbol{\pi}^{\prime})=t_{2}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime})$. Then,
@@
[…]
$=\min_{\boldsymbol{\pi}^{\prime\prime}\in\nu_{\mathcal{F}}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})}\frac{1}{2\sigma^{2}}\frac{\|\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}^{\prime\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}^{\prime\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\min_{\boldsymbol{l}\in\mathcal{L}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}})}\left\{\frac{1}{2\sigma^{2}}\frac{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\right\}$
$=\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}})}\frac{1}{2\sigma^{2}}\frac{(t_{1}-t_{2})^{2}\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{(t_{1}-t_{2})^{2}\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\min_{\boldsymbol{l}\in\mathcal{L}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}})}\left\{\frac{1}{2\sigma^{2}}\frac{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\right\}$
$=\min_{\boldsymbol{l}\in\mathcal{L}}\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}$
$\leq\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\mathfrak{s}_{\tilde{A}}$ (12)
@@
Part I. Since $\min_{\boldsymbol{l}\in\mathcal{L}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\geq 0$ by the design of the Lagrangian and the constraints, we get
@@
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\left(1+\mathfrak{s}_{\tilde{A}}\right)\leq\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})$
$\implies$ $T_{\hat{\mathcal{F}},0}^{-1}({\boldsymbol{\mu}})\left(1+\mathfrak{s}_{\tilde{A}}\right)\leq T_{\mathcal{F},0}^{-1}({\boldsymbol{\mu}})$ (14)
@@
$T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\geq\left(1+\mathfrak{s}_{\tilde{A}}\right)\frac{2\sigma^{2}}{\min_{\boldsymbol{\pi}^{\prime\prime}\in\nu_{\mathcal{F}}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})}d_{\boldsymbol{\pi}^{\prime\prime}}^{2}}$
@@ F.2 Impact of Unknown Linear Constraints: Proof of Corollary 1 Part (b)
Part (b) Let $d_{\boldsymbol{\pi}}^{2}\mathrel{\triangleq}\frac{\|\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}\|_{{\boldsymbol{\mu}}{\boldsymbol{\mu}}^{\top}}^{2}}{\|\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi}\|_{2}^{2}}$ be the norm of the projection of $\boldsymbol{\mu}$} on the policy gap $(\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi})$. Then, the characteristic time $T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})$ satisfies $T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\geq\frac{H}{\kappa^{2}}\left(1-\frac{\epsilon}{\gamma}\right)$. $H$ is the sum of squares of gaps. $\kappa$ is the condition numbers of a sub-matrix of $A$ that consists K{K} linearly independent active constraints for $\boldsymbol{\pi}^{\star}$.
@@
Here, $H=\frac{2\sigma^{2}}{\Delta^{2}}$, where $\Delta^{2}=\sum_{a=1}^{K}({\boldsymbol{\mu}}^{*}-{\boldsymbol{\mu}}_{a})^{2}$, i.e the sum of squares of sub-optimal gaps in the arms and $\kappa$ is the condition number of a sub-matrix of $A$ consisting at least K{K} linearly independent active constraints for $\boldsymbol{\pi}^{\star}$. Referring to the proof structure of Corollary 2 in (Carlsson et al.,, 2024) is of independent interest.
@@
[…]